# Provably Efficient Reinforcement Learning with Multinomial Logit Function Approximation

### v1

@@ B.1 Main Proof
We define
@@
where the second inequality holds by $\lVert\phi_{i,h}^{s^{\prime}}\rVert_{2}\leq 1$ and $\widehat{\Sigma}_{i,h}\geq I,\forall i\in[K]$, the last inequality is by the elliptical potential in Lemma 6. This finishes the proof. ∎
@@ B.2.1 Proof of Lemma 7
$\bar{f}_{k,h}(\theta)-\bar{f}_{k,h}(\theta_{h}^{*})$ $=-\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log p_{k,h}^{s^{\prime}}(\theta)+\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\geq 0,$
@@ B.2.2 Proof of Lemma 8
which implies $x_{i}^{\top}Zx_{i}+x_{j}^{\top}Zx_{j}\geq x_{i}^{\top}Zx_{j}+x_{j}^{\top}Zx_{i},\forall x_{i},x_{j}\in\mathbb{R}^{d}$. Let $x_{i}=(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$, we have
@@ B.2.3 Proof of Lemma 9
Notice that $\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$ is a martingale difference sequence. Also, we have
$\left|\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|$ $\leq\left|\left(\bar{g}_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|+\left|\left(g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|$
@@
$\leq\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})^{2}=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}\triangleq A_{k,h}.$
@@
$\leq\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lvert(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})\rvert\leq\sqrt{kU\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}((\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*}))^{2}}\leq 2B.$
@@
where $m=2\log_{2}(kU)$, and the last inequality follows the Bernstein inequality for martingales. Combining above two cases, letting $\tau=\log{\frac{mk^{2}}{\delta/H}}$ and taking the union bound over $k$ and $h\in[H]$, we have with probability at least $1-\delta$, for any $k\in[K],h\in[H]$ it holds that
@@
$=2\sqrt{\gamma_{k}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}}+8B\left(1+\gamma_{k}\right),$
@@
$\sqrt{\gamma_{k}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}}\leq\frac{2\gamma_{k}}{\kappa}+\frac{\kappa}{8}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}$
Then we have
@@ Appendix C Proof of Lemma 2
The gradient of $p_{s,a}(\theta)$ is given by $\nabla p_{s,a}^{s^{\prime}}(\theta)=p_{s,a}^{s^{\prime}}(\theta)\phi_{s,a}^{s^{\prime}}-p_{s,a}^{s^{\prime}}(\theta)\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta)\phi_{s,a}^{s^{\prime\prime}}$. By the mean value theorem, there exists $\bar{\theta}=\nu\theta_{h}^{*}+(1-\nu)\widehat{\theta}_{k,h}$ for some $\nu\in[0,1]$, such that
@@ Appendix E Proof of Lemma 3
Then, we present the following lemma.
@@
For the update rule defined in (27) as
$\widetilde{\theta}_{k+1,h}=\operatorname*{arg\,min}_{\theta\in\Theta}\widetilde{f}_{k,h}(\theta)+\frac{1}{2\eta}\lVert\theta-\widetilde{\theta}_{k,h}\rVert_{\mathcal{H}_{k,h}},$
let $\eta=\frac{1}{2}\log(U+1)+(B+1)$, we have
@@
Then, we bound the right-hand side of the above lemma separately. First, we bound the first term. Define the softmax function as follows.
@@
$\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})=\underbrace{\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})}_{(a)}+\underbrace{\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})}_{(b)}$
@@
Thus, by setting $\eta=\frac{1}{2}\log(U+1)+(B+1)$, $\lambda=84\sqrt{2}\eta(B+d)$, we have $\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}\leq\mathcal{O}\big{(}\sqrt{d}\log U\log(KH/\delta)\big{)}\triangleq\widetilde{\beta}_{k}$. This finishes the proof. ∎
Appendix F Proof of Theorem 2
@@
By the same analysis as the proof of Lemma 2 in Appendix C, we have
$\bigg{|}{\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}\bigg{|}\leq$ $H\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\phi_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h}-\theta_{h}^{*})$
[…]
where the first inequality holds by (24), the second inequality is by Holder’s inequality, and the third inequality is by Lemma 3.
Then, substituting $\Gamma_{h,s,a}=H\widetilde{\beta}_{k}\max_{s^{\prime}\in\mathcal{S}_{h,s,a}}\lVert\phi_{s,a}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}$ into Lemma 10, we have
$\sum_{k=1}^{K}V_{1}^{*}(s_{k,1})-\sum_{k=1}^{K}V_{1}^{\pi_{k}}(s_{k,1})\leq$ $2H\widetilde{\beta}_{k}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}+H\sqrt{2KH\log(2/\delta)}$
[…]
where the second inequality holds by the Cauchy-Schwarz inequality, and the third inequality is by Lemma 6. This finishes the proof. ∎
Appendix G Proof of Lemma 4
Proof.
By the second-order Taylor expansion, there exists $\bar{\theta}=\nu\theta_{h}^{*}+(1-\nu)\widetilde{\theta}_{k,h}$ for some $\nu\in[0,1]$, such that
${\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widetilde{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}$
@@ Appendix G Proof of Lemma 4
The gradient of $p_{s,a}(\theta)$ is given by $\nabla p_{s,a}^{s^{\prime}}(\theta)=p_{s,a}^{s^{\prime}}(\theta)\phi_{s,a}^{s^{\prime}}-p_{s,a}^{s^{\prime}}(\theta)\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta)\phi_{s,a}^{s^{\prime\prime}}$. For the first-order term, we have
@@
where the inequality holds by $V(s)\in[0,H],\forall s$. According to Lemma 18, we have (omit the subscript $\mathcal{S}_{h,s,a}$ for simplicity):
@@
where the last inequality holds by Lemma 3.
Combining (28) and (31) finishes the proof. ∎
Appendix H Proof of Theorem 3
@@
$\epsilon_{k,h}^{\mathtt{fst}}=H\widetilde{\beta}_{k}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}},\epsilon_{k,h}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}.$
@@
Bounding $\epsilon_{k,h}^{\mathtt{fst}}$.
For simplicity, we denote
@@
$+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(e)}$
Next, we bound these terms separately.
@@
where the fifth inequality is by Cauchy-Schwarz inequality and the fourth inequality is because by Lemma 17, since ~$\widetilde{\mathcal{H}}_{k,h}\succeq\mathcal{H}_{k,h}\succeq\lambda I_{d}$, we have
@@
$\sum_{k=1}^{K}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\leq$ $\sqrt{\sum_{k=1}^{K}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})}\sqrt{\sum_{k=1}^{K}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}^{2}}$
@@
where the last inequality holds by Lemma 16.
@@
$\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{snd}}=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}\leq\frac{5}{\kappa}H^{2}\widetilde{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda}\right).$ (36)
@@
Appendix I Proof of Theorem 4

### v2

@@ B.1 Main Proof
The main proof follows the proof of Theorem 1 in Zhang et al. [2016]. We define
@@
where the second inequality holds by $\lVert\phi_{i,h}^{s^{\prime}}\rVert_{2}\leq 1$ and $\widehat{\Sigma}_{i,h}\geq I,\forall i\in[K]$, and the last inequality is by Lemma 6. This finishes the proof. ∎
@@ B.2.1 Proof of Lemma 7
$\bar{f}_{k,h}(\theta)-\bar{f}_{k,h}(\theta_{h}^{*})$ $=\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\theta_{h}^{*})\log\frac{p_{k,h}^{s^{\prime}}(\theta_{h}^{*})}{p_{k,h}^{s^{\prime}}(\theta)}\geq 0,$
@@ B.2.2 Proof of Lemma 8
which implies $x_{i}^{\top}Zx_{i}+x_{j}^{\top}Zx_{j}\geq x_{i}^{\top}Zx_{j}+x_{j}^{\top}Zx_{i},\quad\forall x_{i},x_{j}\in\mathbb{R}^{d}$.
Let $x_{i}=(p_{k,h}^{s^{\prime}}(\theta)-y_{k,h}^{s^{\prime}})\phi_{k,h}^{s^{\prime}}$, we have
@@ B.2.3 Proof of Lemma 9
First, notice that $\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})$ is a martingale difference sequence. Also, we have
$\quad\left|\left(\bar{g}_{i,h}(\theta_{i})-g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|$
$\leq\left|\left(\bar{g}_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|+\left|\left(g_{i,h}(\theta_{i})\right)^{\top}(\theta_{i}-\theta^{*})\right|$
@@
$\leq\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})^{2}$
$=\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}$
$\triangleq A_{k,h}.$
@@
$\leq\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}\lvert(\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*})\rvert$
$\leq\sqrt{kU\sum_{i=1}^{k}\sum_{s^{\prime}\in\mathcal{S}_{i,h}}((\phi_{i,h}^{s^{\prime}})^{\top}(\theta_{i}-\theta^{*}))^{2}}$
$\leq 2B.$
where the second equality is due to the definition of $\bar{g}_{i,h}(\theta_{i})$ and $g_{i,h}(\theta_{i})$, the first inequality holds by $p_{i,h}^{s^{\prime}}(\theta_{i})-y_{i,h}^{s^{\prime}}\in[-1,1]$, the second inequality holds by the Cauchy-Schwarz inequality, and the last inequality holds by the condition of case (i).
@@
[…]
where $m=2\log_{2}(kU)$, and the last inequality follows the Bernstein inequality for martingales.
Combining above two cases, letting $\tau=\log{\frac{mk^{2}}{\delta/H}}$ and taking the union bound over $k$ and $h\in[H]$, we have with probability at least $1-\delta$, for any $k\in[K],h\in[H]$ it holds that
@@
$=2\sqrt{\gamma_{k}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}}+8B\left(1+\gamma_{k}\right),~{}$ (23)
@@
$\sqrt{\gamma_{k}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}}\leq\frac{2\gamma_{k}}{\kappa}+\frac{\kappa}{8}\sum_{i=1}^{k}\lVert\theta_{i}-\theta^{*}\rVert_{W_{i,h}}^{2}.$ (24)
Combining (23) and (24), we have
@@ Appendix C Proof of Lemma 2
The gradient of $p_{s,a}(\theta)$ is given by
$\nabla p_{s,a}^{s^{\prime}}(\theta)=p_{s,a}^{s^{\prime}}(\theta)\phi_{s,a}^{s^{\prime}}-p_{s,a}^{s^{\prime}}(\theta)\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta)\phi_{s,a}^{s^{\prime\prime}}.$
By the mean value theorem, there exists $\bar{\theta}=\nu\theta_{h}^{*}+(1-\nu)\widehat{\theta}_{k,h}$ for some $\nu\in[0,1]$, such that
@@
=\ = $\bigg{|}\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}\nabla p_{s,a}^{s^{\prime}}(\bar{\theta})(\widehat{\theta}_{k,h}-\theta_{h}^{*})V(s^{\prime})\bigg{|}$
@@ Appendix E Proof of Lemma 3
Then, we present the following lemma that bounds the estimation error of the update rule in (28).
@@
For the update rule defined in (28), it holds that
@@
We bound the right-hand side of the above lemma separately in the following.
First, we bound the first term. Define the softmax function as follows.
@@
$\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})$
=\ = $\underbrace{\sum_{i=1}^{k}f_{i,h}(\theta_{h}^{*})-\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})}_{(a)}+\underbrace{\sum_{i=1}^{k}f_{i,h}(z_{i,h},y_{i,h})-\sum_{i=1}^{k}f_{i,h}(\widetilde{\theta}_{i+1,h})}_{(b)}$
@@
[…]
Thus, by setting $\eta=\frac{1}{2}\log(U+1)+(B+1)$, $\lambda=84\sqrt{2}\eta(B+d)$, we have
$\lVert\widetilde{\theta}_{k+1,h}-\theta_{h}^{*}\rVert_{\mathcal{H}_{k+1,h}}\leq\mathcal{O}\big{(}\sqrt{d}\log U\log(KH/\delta)\big{)}\triangleq\widetilde{\beta}_{k}.$
This finishes the proof. ∎
Appendix F Proof of Lemma 4
@@
By the second-order Taylor expansion, there exists $\bar{\theta}=\nu\theta_{h}^{*}+(1-\nu)\widehat{\theta}_{k,h}$ for some $\nu\in[0,1]$, such that
${\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\widehat{\theta}_{k,h})V(s^{\prime})-\sum_{s^{\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime}}(\theta_{h}^{*})V(s^{\prime})}$
@@
The gradient of $p_{s,a}(\theta)$ is given by
$\nabla p_{s,a}^{s^{\prime}}(\theta)=p_{s,a}^{s^{\prime}}(\theta)\phi_{s,a}^{s^{\prime}}-p_{s,a}^{s^{\prime}}(\theta)\sum_{s^{\prime\prime}\in\mathcal{S}_{h,s,a}}p_{s,a}^{s^{\prime\prime}}(\theta)\phi_{s,a}^{s^{\prime\prime}}.$
For the first-order term, we have
@@
where the inequality holds by $V(s)\in[0,H],\forall s$.
According to Lemma 18, we have (omit the subscript $\mathcal{S}_{h,s,a}$ for simplicity):
@@
where the last inequality holds by Lemma 3. Combining (29) and (32) finishes the proof. ∎
Appendix G Proof of Theorem 2
@@
$\epsilon_{k,h}^{\mathtt{fst}}$ $=H\widetilde{\beta}_{k}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k,h})\Big{\lVert}\phi_{k,h}^{s^{\prime}}-\sum_{s^{\prime\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime\prime}}(\widetilde{\theta}_{k,h})\phi_{k,h}^{s^{\prime\prime}}\Big{\rVert}_{\mathcal{H}_{k,h}^{-1}},$
$\epsilon_{k,h}^{\mathtt{snd}}$ $=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}.$
@@
Bounding $\epsilon_{k,h}^{\mathtt{fst}}$.  For simplicity, we denote
@@
$+\underbrace{\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}}_{(e)}.$
We bound these terms separately in the following.
@@
[…]
where and the fifth inequality is by Cauchy-Schwarz inequality and the fourth inequality is because by Lemma 17, since ~$\widetilde{\mathcal{H}}_{k,h}\succeq\mathcal{H}_{k,h}\succeq\lambda I_{d}$, we have
@@
$\hskip 56.9055pt+\frac{2\eta}{\sqrt{\lambda}}\frac{\left(\max_{s^{\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\right)^{2}+\left(\max_{s^{\prime\prime\prime}\in\mathcal{S}_{k,h}}\big{\lVert}\phi_{k,h}^{s^{\prime\prime\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}\right)^{2}}{2}$
@@
$\sum_{k=1}^{K}\sum_{s^{\prime}\in\mathcal{S}_{k,h}}p_{k,h}^{s^{\prime}}(\widetilde{\theta}_{k+1,h})\big{\lVert}\widetilde{\phi}_{k,h}^{s^{\prime}}\big{\rVert}_{\mathcal{H}_{k,h}^{-1}}$
@@
[…]
where the first inequality holds by the Cauchy-Schwarz inequality and the last holds by Lemma 6.
@@
$\sum_{k=1}^{K}\sum_{h=1}^{H}\epsilon_{k,h}^{\mathtt{snd}}$ $=\frac{5}{2}H\widetilde{\beta}_{k}^{2}\sum_{k=1}^{K}\sum_{h=1}^{H}\max_{s^{\prime}\in\mathcal{S}_{k,h}}\lVert\phi_{k,h}^{s^{\prime}}\rVert_{\mathcal{H}_{k,h}^{-1}}^{2}$
$\leq\frac{5}{\kappa}H^{2}\widetilde{\beta}_{K}^{2}d\log\left(1+\frac{K}{d\lambda}\right).$ (37)
where the last inequality holds by Lemma 6.
@@
Appendix H Proof of Theorem 3